OpenStudy (gabbyalicorn):

The width of a rectangle is 6 in. less than its length. The perimeter is 68 in. What is the width of the rectangle? in.

OpenStudy (gabbyalicorn):

Hi pooja.

pooja195 (pooja195):

Uh i think the answer is in the question....if im right

OpenStudy (anonymous):

divide it in two first

OpenStudy (jhannybean):



OpenStudy (jhannybean):

Omg I forgot the shapes have the "fill" option.... anyway.

OpenStudy (anonymous):

yeah

OpenStudy (gabbyalicorn):

When you say divide two first, do you mean 68 ÷ 2 and 6÷ 2

OpenStudy (anonymous):

no 68 by 2

OpenStudy (anonymous):

then subtract 6 from that

OpenStudy (gabbyalicorn):

okay, that is 34...

OpenStudy (jhannybean):

the length of the rectangle is given, but they say the `width` `is (equal to)` the `length - 6 in`You can rewrite this as \(\sf w = L-6\)

OpenStudy (jhannybean):

So the Perimeter of a rectangle is represented as: \[\sf P = 2L +2W\]

OpenStudy (jhannybean):

Since we're given the perimeter, and that is = 68 in. we can substitute that in to our formula. \[\sf 68 = 2L +2W\]

OpenStudy (jhannybean):

But we also know something else, \(\sf W=L-6\)

OpenStudy (gabbyalicorn):

68 - 6?

OpenStudy (jhannybean):

So we can substitute this in to the formula for the perimeter. \[\sf 68 = 2L+2(L-6)\]

OpenStudy (jhannybean):

Are you following, @gabbyalicorn ?

OpenStudy (gabbyalicorn):

Um, somewhat... I don't really get it but i'm following the steps if that's what you mean. :)

OpenStudy (jhannybean):

Which part are you confused with? Let's clarify that before moving on.

OpenStudy (gabbyalicorn):

is L the variable we are trying to find

OpenStudy (gabbyalicorn):

the number for...

OpenStudy (jhannybean):

We are trying to find W (width). The problem indiscreetly gives us the equation for the width by saying that the width IS 6 LESS THAN the length. We can write an equation for that. \(\sf w = L-6\)

OpenStudy (gabbyalicorn):

but what is L

OpenStudy (jhannybean):

But we can't automatically find the width, \(\sf w\) can we? We need to find the length, \(\sf L\) first.

OpenStudy (jhannybean):

That is where we use the equation for the perimeter of the rectangle. \(\sf P = 2L +2w\)

OpenStudy (jhannybean):

Do you see how this works?

OpenStudy (gabbyalicorn):

I think I'm getting it...

OpenStudy (jhannybean):

Okay, so we have \[\sf 68 = 2L+2(L-6)\] We need to first distribute (meaning multiply) 2 to each term inside the parenthesis. \[2 \cdot L =~?\]\[2\cdot (-6) =~?\]

OpenStudy (gabbyalicorn):